# brick-splitting Torsion Pairs and Trim Lattices

### v1

@@ 4.2. Construction of new brick-directed algebras
Let $A=kQ/I$ and $B=kQ^{\prime}/I^{\prime}$ be two connected algebras, where $I$ and $I^{\prime}$ are assumed to be admissible ideals. Let $R$ and $R^{\prime}$ denote two sets of admissible relations, respectively, in $kQ$ and $kQ^{\prime}$, which generate $I$ and $I^{\prime}$. Moreover, assume that $Q_{0}=\{x_{1},\ldots,x_{n}\}$ and $Q^{\prime}_{0}=\{y_{1},\ldots,y_{m}\}$ denote, respectively, the vertex set of $A$ and $B$. Provided that $x_{n}$ is a sink in $Q$, and $y_{1}$ is a source in $Q^{\prime}$, then by $\Lambda:=k\tilde{Q}/\tilde{I}$ we denote the algebra obtained from gluing of $A$ and $B$ at the vertices $x_{n}$ and $y_{1}$, with the bound quiver $(\tilde{Q},\tilde{I})$ constructed as follows:
@@
From the construction, it is immediate that $(\tilde{Q},\tilde{I})$ contains a copy of $(Q,I)$, as well as a copy of $(Q^{\prime},I^{\prime})$, as bound subquivers. Moreover, other than vertex $v$ in $\tilde{Q}$, any other vertex in $\tilde{Q}_{0}$ belongs either to $Q_{0}$ or $Q^{\prime}_{0}$, and not both. We note that because $B$ has a source, it is immediate that $\Lambda=k\tilde{Q}/\tilde{I}$ is isomorphic to $A=kQ/I$ if and only if the bound quiver of $B$ consists of only one vertex, that is, if $B\simeq k$.
@@
Let $A=kQ/I$ and $B=kQ^{\prime}/I^{\prime}$ be two connected algebras. If $\Lambda=k\tilde{Q}/\tilde{I}$ is obtained from gluing $A$ and $B$, then $\Lambda$ is a connected algebra of rank $|A|+|B|-1$. Furthermore, the following hold:
@@
where all arrows $\beta_{i}$, for $11\leq i\leq n-$ 1, are equi-oriented, that is, $\beta_{i}:i\rightarrow i+1$. In particular, the bar $\mathbf{b}:=\beta_{n-1}\cdots\beta_{1}$, which starts at $1$ and ends at $n$, is a copy of the equi-oriented Dynkin quiver $A_{n}$.
@@
We also use these concrete examples to highlight another fundamental difference between representation-directed algebras and brick-directed algebras. First, we recall that if $A$ is a connected representation-finite algebra, then $A$ is representation-directed if and only if the (unique) component of $\Gamma_{A}$ is preprojective (see [ASS, IX.3, Lemma 3.5]). That is, among connected algebras, representation-directed algebras can be classified as those whose Auslander-Reiten quiver consists of a unique preprojective component. However, brick-directed algebras may or may not admit a preprojective component. For instance, the Kronecker algebra is brick-directed (see Prop. 4.6) and it is known to admit a (unique infinite) preprojective component. In contrast, the wind-wheel algebras in Example 4.14, which are shown to be brick-directed, do not admit any preprojective components. This is because they are representation-infinite but brick-finite (also see [Mo1]). Also, Example 3.5 gives an explicit example of a connected representation-finite algebra which is brick-directed without a preprojective component.
@@
For (IV), let $A$ be the wild local algebra $k[x_{1},x_{2},x_{3}]/\langle x_{i}x_{j}:1\leq i\leq j\leq 3\rangle$, whose bound quiver $(Q,I)$ consists of a single vertex and $3$ loops, say $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, subject to all quadratic relations. Since $\operatorname*{brick}A$ consists only of the unique simple $A$-module, $A$ is evidently brick-finite and brick-directed. Moreover, let $Q^{\prime}$ be the equi-oriented Dynkin quiver $1\xrightarrow{\beta_{1}}2\xrightarrow{\beta_{2}}\cdots n-1\xrightarrow{\beta_{n-1}}n$. As remarked previously, $B=kQ^{\prime}$ is a (representation-finite) brick-directed algebra. Now, we construct the desired algebra $\Lambda=k\tilde{Q}/\tilde{I}$ from $A$ and $B$, similar to the gluing construction described before Proposition 4.13. More precisely, identify vertex $n$ in $Q^{\prime}$ with the unique vertex of $Q$ to obtain the new quiver $\tilde{Q}$, and let $\tilde{I}:=\langle\alpha_{i}\alpha_{j},\alpha_{i}\beta_{n-1}:1\leq i\leq j\leq 3\rangle$. Observe that $A$ is a proper quotient of $\Lambda$, hence $\Lambda$ is wild. Observe that if $X$ is a brick over $\Lambda$, then either $X(\beta_{n-1})=0$ or $X(\alpha_{i})=0$ for all $1\leq i\leq 3$. Otherwise, the simple module $S_{n}$ appears both in the socle and the top of $X$, so $X$ cannot be a brick. Consequently, $\operatorname*{brick}\Lambda=\operatorname*{brick}B$, so $\Lambda$ is brick-finite and brick-directed.
@@ 4.3. Wall-chamber structure of brick-directed algebras
[…]
@@
$\overline{\mathcal{T}}_{\theta}$ :=$:=\{M\in\operatorname*{mod}A\mid\text{$\theta([N])\geq 0$ for every quotient module $N$ of $M$}\},$
$\mathcal{F}_{\theta}$ :=$:=\{M\in\operatorname*{mod}A\mid\text{$\theta([L])<0$ for every submodule $L\neq 0$ of $M$}\},$
$\mathcal{T}_{\theta}$ :=$:=\{M\in\operatorname*{mod}A\mid\text{$\theta([N])>0$ for every quotient module $N\neq 0$ of $M$}\},$
¯$\overline{\mathcal{F}}_{\theta}$ :=$:=\{M\in\operatorname*{mod}A\mid\text{$\theta([L])\geq 0$ for every submodule $L$ of $M$}\}.$
@@
5. Trim lattices in representation theory
@@
In what follows, we first prove Theorem 1.7. This is obtained as a consequence of the next two propositions, which can be seen as the algebraic counterparts of Theorems 2.1 and 2.2, where we are concerned with the lattice of torsion classes of brick-finite algebras.
@@
By Corollary 5.3, $A$ is brick-directed if and only if $\operatorname*{tors}A$ is a trim lattice. Moreover, by Proposition 4.4, the algebras $A/J$, $eAe$, and $B$ described above are also brick-directed. Hence, the implications follow immediately from Propositions 5.1 and 5.2. ∎
@@
Let $A$ be a brick-finite algebra. If $A$ is brick-directed, then each brick in $\operatorname*{mod}A$ is uniquely determined by its dimension vector, that is, for two distinct elements $X$ and $Y$ in $\operatorname*{brick}A$, we have $[X]\neq[Y]$.
@@
For any two elements $v,w\in K_{0}(\operatorname*{mod}A)_{\mathbb{R}}$, we write $v\leq w$ if $w-v=\sum_{i=1}^{n}a_{i}[S_{i}]$ with $a_{i}\geq 0$ for each $i$, and $v<w$ means $v\leq w$ and $v\neq w$. We call a sequence $v_{0}<v_{1}<\cdots<v_{l}$ an increasing path of the Newton polytope $\mathrm{N}(M)$ if $v_{i}$ is a vertex of $\mathrm{N}(M)$, for each $i\in\{0,\ldots,l\}$, and the line segment $[v_{i-1},v_{i}]$ is an edge of $\mathrm{N}(M)$ with $v_{i-1}<v_{i}$, for each $i\in\{1,\ldots,l\}$.
@@
    {maximal chains in torsA}{increasing paths in N(M) from 0 to [$\{\text{maximal chains in $\operatorname*{tors}A$}\}\to\{\text{increasing paths in $\mathrm{N}(M)$ from $0$ to $[M]$}\},$
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of $X\in\operatorname*{brick}A$. Assume that $0=v_{0}<v_{1}<\cdots<v_{l}=[M]$ is an increasing path in $\mathrm{N}(M)$ from 00 to $[M]$. Then the path is indivisible if and only if $l=|{\operatorname*{brick}A}|$.
@@
We first show the “if” part. The bricks $X_{1},\ldots,X_{l}$ are distinct by Proposition 2.6, so the assumption $l=|{\operatorname*{brick}A}|$ implies that $X_{1},\ldots,X_{l}$ are all the bricks in $\operatorname*{mod}A$. Thus $\sum_{X\in\operatorname*{brick}A}[X]=\sum_{i=1}^{l}m_{i}[X_{i}]$ above gives $m_{i}=1$, for any $i$. This implies that the increasing path $0=v_{0}<v_{1}<\cdots<v_{l}=[M]$ is indivisible, because Lemma 4.18(2) implies that each $[X_{i}]$ is indivisible.
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of $X\in\operatorname*{brick}A$. Then, $A$ is brick directed if and only if there exists an indivisible increasing path in $\mathrm{N}(M)$ from 00 to $[M]$.
@@
As in the proof of Proposition 5.2, $A$ is brick directed if and only if $\operatorname*{tors}A$ has a maximal chain of length $|{\operatorname*{brick}A}|$. The latter is equivalent to the existence of an increasing path of $\mathrm{N}(M)$ of length $|{\operatorname*{brick}A}|$ from 00 to $[M]$, by Proposition 5.9(2), and also to the existence of an indivisible increasing path of $\mathrm{N}(M)$ from 00 to $[M]$, by Lemma 5.10. ∎
@@
[…]
@@
    As observed previously, there exist various families of brick-infinite algebras that are brick-directed (see Prop. 4.6 and Cor. 4.16). We note that, in contrast to the uniqueness property in Corollary 5.7, one cannot expect a similar statement for the brick-infinite brick-directed algebras. More precisely, let $A$ be a brick-infinite algebra over an algebraically closed field. Then, by the so-called “Second brick-Brauer-Thrall” conjecture (2nd bBT), first posed by the third-named author in 2019, it is expected that $\operatorname*{brick}A$ contains an infinite set of bricks with the same dimension vector. For the original motivations of the 2nd bBT, and the most recent state of art on this conjecture, respectively, see [Mo2, Mo3], [MP2, MP3], and the references therein. Meanwhile, one can naturally ask if, for an arbitrary brick-directed algebra $A$, and for each dimension vector $\underline{d}\in\mathbb{Z}^{n}_{\geq 0}$, the module variety $\operatorname*{mod}(A,\underline{d})$ always admits at most one brick component (for terminology and some motivations, see [MP1]).
Acknowledgments. We thank Hugh Thomas for some stimulating discussions about trim lattices at the earlier stage of this project, and for some helpful follow-up comments. We also thank Adrien Segovia for drawing our attention to Theorem 2.2. SA was supported by Early-Career Scientist JSPS Kakenhi grant number 23K12957. OI was supported by JSPS Grant-in-Aid for Scientific Research (B) 22H01113, (B) 23K22384. KM was supported by Early-Career Scientist JSPS Kakenhi grant number 24K16908. CP was supported by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2018-04513) and by the Canadian Defence Academy Research Programme.

### v2

@@ 4.3. Construction of new brick-directed algebras
[…]
@@
Let $A=kQ/I$ and $B=kQ^{\prime}/I^{\prime}$ be two connected algebras. If $\Lambda=k\tilde{Q}/\tilde{I}$ is obtained from gluing $A$ and $B$, then $\Lambda$ is a connected algebra of rank |A|+|B|1|A|+|B|-1. Furthermore, the following hold:
@@
where all arrows $\beta_{i}$, for $1\leq i\leq n-1$, are equi-oriented, that is, $\beta_{i}:i\rightarrow i+1$. In particular, the bar $\mathbf{b}:=\beta_{n-1}\cdots\beta_{1}$, which starts at $1$ and ends at $n$, is a copy of the equi-oriented Dynkin quiver $A_{n}$.
@@
We also use these concrete examples to highlight another fundamental difference between representation-directed algebras and brick-directed algebras. First, we recall that if $A$ is a connected representation-finite algebra, then $A$ is representation-directed if and only if the (unique) component of $\Gamma_{A}$ is preprojective (see [ASS, IX.3, Lemma 3.5]). That is, among connected algebras, representation-directed algebras can be classified as those whose Auslander-Reiten quiver consists of a unique preprojective component. However, brick-directed algebras may or may not admit a preprojective component. For instance, the Kronecker algebra is brick-directed (see Prop. 4.9) and it is known to admit a (unique infinite) preprojective component. In contrast, the wind-wheel algebras in Example 4.11, which are shown to be brick-directed, do not admit any preprojective components. This is because they are representation-infinite but brick-finite (also see [Mo1]). Also, Example 3.12 gives an explicit example of a connected representation-finite algebra which is brick-directed without a preprojective component.
@@
For (IV), let $A$ be the wild local algebra $k[x_{1},x_{2},x_{3}]/\langle x_{i}x_{j}:1\leq i\leq j\leq 3\rangle$, whose bound quiver $(Q,I)$ consists of a single vertex and $3$ loops, say $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, subject to all quadratic relations. Since $\operatorname*{brick}A$ consists only of the unique simple $A$-module, AA is evidently brick-finite and brick-directed. Moreover, let $Q^{\prime}$ be the equi-oriented Dynkin quiver $1\xrightarrow{\beta_{1}}2\xrightarrow{\beta_{2}}\cdots n-1\xrightarrow{\beta_{n-1}}n$. As remarked previously, $B=kQ^{\prime}$ is a (representation-finite) brick-directed algebra. Now, we construct the desired algebra $\Lambda=k\tilde{Q}/\tilde{I}$ from $A$ and $B$, similar to the gluing construction described before Proposition 4.10. More precisely, identify vertex $n$ in $Q^{\prime}$ with the unique vertex of $Q$ to obtain the new quiver $\tilde{Q}$, and let $\tilde{I}:=\langle\alpha_{i}\alpha_{j},\alpha_{i}\beta_{n-1}:1\leq i\leq j\leq 3\rangle$. Observe that $A$ is a proper quotient of $\Lambda$, hence $\Lambda$ is wild. Observe that if $X$ is a brick over $\Lambda$, then either $X(\beta_{n-1})=0$ or $X(\alpha_{i})=0$ for all $1\leq i\leq 3$. Otherwise, the simple module $S_{n}$ appears both in the socle and the top of $X$, so $X$ cannot be a brick. Consequently, $\operatorname*{brick}\Lambda=\operatorname*{brick}B$, so $\Lambda$ is brick-finite and brick-directed.
@@ 4.4. Wall-chamber structure of brick-directed algebras
$\Theta_{M}:=\{\theta\in K_{0}(\operatorname*{proj}A)_{\mathbb{R}}\mid\text{$M$ is $\theta$-semistable}\}.$
@@
[…]
@@
5. Left modular lattices in representation theory
@@
In what follows, we first prove Theorem 1.7. This is obtained as a consequence of the next two propositions, which can be seen as the algebraic counterparts of Theorem 2.1, where we are concerned with the lattice of torsion classes of brick-finite algebras.
@@
By Corollary 5.3, AA is brick-directed if and only if $\operatorname*{tors}A$ is a trim lattice. Moreover, by Proposition 4.7, the algebras $A/J$, eAeeAe, and $B$ described above are also brick-directed. Hence, the implications follow immediately from Propositions 5.1 and 5.2. ∎
@@
Let $A$ be a brick-finite algebra. If $A$ is brick-directed, then each brick in $\operatorname*{mod}A$ is uniquely determined by its dimension vector, that is, for two distinct elements $X$ and $Y$ in $\operatorname*{brick}A$, we have $X]\neq[Y]$.
@@
For any two elements $v,w\in K_{0}(\operatorname*{mod}A)_{\mathbb{R}}$, we write $v\leq w$ if $w-v=\sum_{i=1}^{n}a_{i}[S_{i}]$ with $a_{i}\geq 0$ for each $i$, and $v<w$ means $v\leq w$ and $v\neq w$. We call a sequence $v_{0}<v_{1}<\cdots<v_{l}$ an increasing path of the Newton polytope $\mathrm{N}(M)$ if $v_{i}$ is a vertex of $\mathrm{N}(M)$, for each $i\in\{0,\ldots,l\}$, and the line segment $v_{i-1},v_{i}]$ is an edge of $\mathrm{N}(M)$ with $v_{i-1}<v_{i}$, for each $i\in\{1,\ldots,l\}$.
@@
[…]
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of all $X\in\operatorname*{brick}A$. Assume that $0=v_{0}<v_{1}<\cdots<v_{l}=[M]$ is an increasing path in $\mathrm{N}(M)$ from 0 to $[M]$. Then the path is indivisible if and only if $l=|{\operatorname*{brick}A}|$.
@@
We first show the “if” part. The bricks $X_{1},\ldots,X_{l}$ are distinct by Proposition 2.5, so the assumption $l=|{\operatorname*{brick}A}|$ implies that $X_{1},\ldots,X_{l}$ are all the bricks in $\operatorname*{mod}A$. Thus $\sum_{X\in\operatorname*{brick}A}[X]=\sum_{i=1}^{l}m_{i}[X_{i}]$ above gives $m_{i}=1$, for any $i$. This implies that the increasing path $0=v_{0}<v_{1}<\cdots<v_{l}=[M]$ is indivisible, because Lemma 4.15(2) implies that each $X_{i}]$ is indivisible.
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of all $X\in\operatorname*{brick}A$. Then, $A$ is brick directed if and only if there exists an indivisible increasing path in $\mathrm{N}(M)$ from 0 to $[M]$.
@@
As in the proof of Proposition 5.2, $A$ is brick directed if and only if $\operatorname*{tors}A$ has a maximal chain of length $|{\operatorname*{brick}A}|$. The latter is equivalent to the existence of an increasing path of $\mathrm{N}(M)$ of length $|{\operatorname*{brick}A}|$ from 0 to $[M]$, by Proposition 5.9(2), and also to the existence of an indivisible increasing path of $\mathrm{N}(M)$ from 0 to $[M]$, by Lemma 5.10. ∎
@@
    ${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$
@@
    As observed previously, there exist various families of brick-infinite algebras that are brick-directed (see Prop. 4.9 and Cor. 4.13). We note that, in contrast to the uniqueness property in Corollary 5.7, one cannot expect a similar statement for the brick-infinite brick-directed algebras. More precisely, let $A$ be a brick-infinite algebra over an algebraically closed field. Then, by the so-called “Second brick-Brauer-Thrall” conjecture (2nd bBT), first posed by the third-named author in 2019, it is expected that $\operatorname*{brick}A$ contains an infinite set of bricks with the same dimension vector. For the original motivations of the 2nd bBT, and the most recent state of art on this conjecture, respectively, see [Mo2, Mo3], [MP2, MP3], and the references therein. Meanwhile, one can naturally ask if, for an arbitrary brick-directed algebra $A$, and for each dimension vector $\underline{d}\in\mathbb{Z}^{n}_{\geq 0}$, the module variety $\operatorname*{mod}(A,\underline{d})$ always admits at most one brick component (for terminology and some motivations, see [MP1, MP4]).
Acknowledgments. We thank Hugh Thomas for some stimulating discussions about trim lattices at the earlier stage of this project, and for some helpful follow-up comments. We also thank Adrien Segovia for drawing our attention to [Mu, Theorem 3.2], and Eric Hanson for pointing out some related studies in [BH, Section 5]. SA was supported by Early-Career Scientist JSPS Kakenhi grant number 23K12957. OI was supported by JSPS Grant-in-Aid for Scientific Research (B) 22H01113, (B) 23K22384. KM was supported by Early-Career Scientist JSPS Kakenhi grant number 24K16908. CP was supported by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2018-04513) and by the Canadian Defence Academy Research Programme.